Point charge 2.5
μC
μC
is located at
x
x
= 0,
y
y
= 0.30
m
m
, point charge -2.5
μC
μC
is located at
x
x
= 0
y
y
= -0.30
m
m
. What are (a)the magnitude and (b)direction of the total electric force that these charges exert on a third point charge
Q
QQ
= 5.0
μC
μC
at
x
x
= 0.40
m
m
,
y
y
= 0?


Part A
Express your answer using two significant figures.




|F|
|F|
=

nothing
N
N

SubmitRequest Answer
Part B




θ
θ
=

nothing



clockwise from the
+x
+x
direction
1
Expert's answer
2020-01-21T03:32:41-0500


the above picture represent the actual case of the above problem where F1 is the force exerted by the charge q1 and F2 is the force exerted by the q2 and F1 and F2 are equal in magnitude and F(resultant) is the net force of F1 and F2

force is given by formula "F=Kq_1q_2\/r^2"

F1=F2= 9*109*2.5*10-6*5*10-6/(0.5)2

F1=F2= 0.45 N

F(resultant)=2*0.45*cos"\theta" = 2*0.45*3/5 = 0.54 N